OpenStudy (ememnim):

Algebra:

OpenStudy (ememnim):

Perform the indicated operation and express the result as a simplified complex number. (−4 + 5i) − (2 − i)

satellite73 (satellite73):

step one \[-4+5i-2+i\] step two is combine like terms

OpenStudy (ememnim):

Perform the indicated operation and express the result as a simplified complex number. (1 + 4i)(7 − i)

satellite73 (satellite73):

\[(a+bi)(c+di)=ac-bd+(ad+bc)i\]

satellite73 (satellite73):

lol \[-\frac{36}{37}-\frac{29}{37}i\] is what it wants

OpenStudy (ememnim):

it still says wrong

satellite73 (satellite73):

lets see

OpenStudy (ememnim):

nevermind dont work on it anymore it wont let me answer again

satellite73 (satellite73):

oh i made a typo, sorry

pooja195 (pooja195):

Woah too many questions! One at a time please.

satellite73 (satellite73):

i guess they want you to multiply like \[\sqrt{2}i(\sqrt{2}i-\sqrt{7}i)\]